1 A renewable energy electricity supply technology has the following characteristics: Capital cost ($) Annual operating cost ($) Lifetime (years) Salvage value ($) Annual electricity supplied (MWh) 300 000 27 200 25 40 000 400 1.1 If the owner can sell the electricity at 25 c/kWh, what is the simple payback period for the technology? 1.2 Would the owner invest in this technology if (s)he set a strict maximum four-year payback period? 1.3 What would the selling price for the electricity have to rise to for the owner to invest in the technology if (s)he set a maximum three-year payback period? 1.4 What is the Present Worth (Net Present Value) of the investment over a 25 year assessment period and real discount rate of 5% when the electricity price is 25 c/kWh? 1.5 What is the real internal rate of return for the owner of this technology over a 25-year assessment period when the electricity price is 30 c/kWh? Would the owner invest if their threshold real rate of return was 30%? 1.6 Derive an analytical relationship between simple payback period and internal rate of return (IRR) over a 15-year assessment period for a project with a single fixed capital payment (K) at the beginning of year 1 and equal constant-dollar annual net benefits over this period (B). Hint: the simple payback period will be K/B. Use the equation for IRR given in the week 2 lecture. Then solve this equation iteratively using Excel for payback periods between 1 and 15, and plot the corresponding graph of IRR vs Payback Period 1.7 With reference to your answers to 1.1 to 1.6, discuss briefly the limitations of the simple payback period as an evaluation criterion and why this can disadvantage renewable energy technologies compared to conventional fossil fuel power supply (at least 200 words). ….Continued on next page 2 2.1 Using the same figures as in question 1, calculate the lifecycle cost of the technology over an assessment period of 25 years at a real discount rate of 5% 2.2 Calculate the average unit cost of the power in present value terms (in cents/kWh) supplied by the technology over its lifetime at this real discount rate. 2.3 What is the corresponding Levelised Cost of Electricity (LCOE) (in cents/kWh)? Why is this value higher than that obtained in question 2.2? (Note: LCOE will not be covered in lectures until week 3.) 2.4 How would the competitiveness of this electricity from a renewable energy source be changed if there was (a) a price on carbon, and (b) a Clean Energy Target? 3 Using the figures in the table in Q1 as a baseline, work out an expression for Present Worth with real discount rate, assessment period, salvage value, and electricity price as independent variables. Then changing just one variable at a time (other things being kept equal) plot graphs of Present Worth versus each of these variables. Use a range of assessment periods up to the lifetime of the technology. Explore the effects of both positive and negative salvage values. On the basis of these graphs and the lectures presented, critically discuss the relative influence of these variables on Present Worth, and hence the more general implications for the economic assessment of renewable energy technologies. You may wish to relate variations in electricity price to carbon pricing. 300 words minimum. Note: to simplify the calculation of present worth, for assessment periods less than the lifetime, neglect the residual value of the technology, and assume salvage values are only incurred at the end of the lifetime of the technology.
Answer #1
Initial cash flow:
Capital cost $300,000
Annual Cash flow:
Annual operating cost $27,200
Number of years(Lifetime) 25
Annual electricity supply (MWh) 400
Terminal Cash flow:
Salvage value $40,000
1.1 CALCULATION OF SIMPLE PAYBACK PERIOD
Selling price per KWh $0.25
A Annual Cash inflow=0.25*400000 $100,000
B Annual Cash out flow(Operating cost) $27,200
C=A-B Net annual Cash inflow $72,800
D=300000/C Simple Payback period(Initial cash flow/72800) 4.120879121 Years
Simple Payback period 4.1 Years
1.2
Under Strict maximum payback period of 4 years;
Owner will not invest in this technology Rejected
1.3 Maximum 3 year payback period:
Annual net cash inflow required=300000/3 $100,000
Annual revenue required=100000+27200 $127,200
Selling price required=127200/400000 $               0.32 per KWh
Selling price required=32 cents/KWh
1.4 CALCULATION OF NET PRESENT VALUE
Net Present value= Sum of (Present Value(PV) of cash flows)
Present Value (PV) of Cash flow=(Cash flow)/((1+i)^N)
i=discount rate=5%=0.05
N=Year of the cash flow
Yearwise cash flow and PV of cash flows are given below:
N A B=A/(1.05^N)
Year Cash flow PV of Cash flow)
0 ($300,000) -300000
1 $72,800 69333.33333
2 $72,800 66031.74603
3 $72,800 62887.37717
4 $72,800 59892.74016
5 $72,800 57040.70492
6 $72,800 54324.48088
7 $72,800 51737.60083
8 $72,800 49273.90556
9 $72,800 46927.5291
10 $72,800 44692.88486
11 $72,800 42564.65225
12 $72,800 40537.76404
13 $72,800 38607.39433
14 $72,800 36768.94698
15 $72,800 35018.04474
16 $72,800 33350.5188
17 $72,800 31762.39886
18 $72,800 30249.90367
19 $72,800 28809.43207
20 $72,800 27437.55435
21 $72,800 26131.00415
22 $72,800 24886.67062
23 $72,800 23701.59106
24 $72,800 22572.94387
(72800+40000 25 $112,800 33310.15265
Total 737851.2753
Net Present value= Sum of (Present Value(PV) of cash flows)= $             737,851
1.5 CALCULATION OF INTERNAL RATE OF RETURN WITH PRICE=30cent/KWh
Annual Sales Revenue with 30 cents per KWh $        120,000
Annual Cash inflow(120000-27200)= $          92,800
Year Cash flow
0 ($300,000)
1 $92,800
2 $92,800
3 $92,800
4 $92,800
5 $92,800
6 $92,800
7 $92,800
8 $92,800
9 $92,800
10 $92,800
11 $92,800
12 $92,800
13 $92,800
14 $92,800
15 $92,800
16 $92,800
17 $92,800
18 $92,800
19 $92,800
20 $92,800
21 $92,800
22 $92,800
23 $92,800
24 $92,800
(92800+40000) 25 $132,800
Internal Rate of Return (IRR) 30.90% Using IRR function of excel over the cash flow)


Book1 Microsoft Excel Home Insert Page LayoutFormulas Data Revi ViewDevelope Cut Copy Paste k ▲ 들를 con attinnalasomat styles